# GenHancer: Imperfect Generative Models are Secretly Strong Vision-Centric Enhancers

### v2

@@ 1 Introduction
Key Point #2: How to design denoising configurations to transfer useful information for visual representations? The structure of the denoiser could determine the enhancement effects. Additionally, before training CLIP, it is essential to mitigate irrelevant information. Therefore, we investigate the influence of different sizes of the denosier and training stages. In this paper, we propose GenHancer, a two-stage post-training method for visual enhancements. In the first stage, we pre-train the projector and denoiser while freezing the ViT, learning basic reconstruction abilities and mitigating irrelevant information. In the second stage, we fine-tune CLIP ViT to enhance its fine-grained visual representations. Meanwhile, we empirically found that a lightweight denoiser is sufficient to achieve remarkable results, which is more efficient yet stronger, as in Fig. 2.
@@ 2 Related Works
Currently, MLLMs predominantly employ CLIP [34] for visual encoding. Tong et al. [42] identified several failure patterns in CLIP, which hinder the fine-grained visual understanding of MLLMs. To overcome this issue, early efforts [42, 41, 18] employed an ensemble of visual experts to combat the visual shortcomings. More recently, ROSS [46] leverages intrinsic visual activations and incorporates a self-supervised visual reconstruction loss during training MLLMs. Complementarily, DIVA [47] proposes to enhance CLIP’s fine-grained abilities through diffusion feedback. Similar to [47], we independently enhance CLIP’s internal representations, which not only strengthens CLIP as a vision-language retriever but also enables the enhanced CLIP to be seamlessly integrated into MLLMs in a plug-and-play manner for better fine-grained vision-centric performance.
@@
Early works [40, 57, 37, 31] utilize generative models as data augmenters. Another line of works [48, 16, 5] leverages self-supervised reconstruction tasks with diffusion models, which helps models grasp visual details and learn fine-grained representations. Similarly, DIVA [47] takes CLIP’s features as conditional inputs to the diffusion model [35], addressing its visual shortcomings through reconstruction. In summary, prior arts predominantly rely on diffusion models [11, 35], whereas we apply our method to both continuous and discrete generative models.
@@
Canonical evaluations of MLLMs focus on fundamental multimodal Q&A capabilities across various domains, e.g., general perception and cognition [10], text and characters [38], scientific fields [30], and potential hallucinations [25, 13] in MLLMs. However, these benchmarks could not effectively assess a model’s fine-grained visual perception abilities, such as object color, quantity, orientation, and viewpoint. To solve this issue, Tong et al. [42] systematically explore the failure modes of CLIP and propose a challenging MMVP benchmark with 9 visual patterns. CV-Bench [41] further expands with 2,600 vision-centric VQA questions, covering dimensions like spatial relationships, count, depth, and distance of both 2D and 3D domains. Besides, NaturalBench [24] curates natural adversarial samples that are easy for humans but MLLMs struggle with. In this paper, we employ these vision-centric benchmarks to comprehensively evaluate models’ fine-grained visual abilities.
@@ 3 Preliminaries of Generative Models
$\mathcal{L}_{\text{CE}}=\mathbb{E}_{\boldsymbol{x}\sim p_{\text{data}}}-\log\prod_{i=1}^{L}\boldsymbol{p}_{\phi}(s_{i}|s_{<i}),\vspace{-7pt}$ (2)
@@ 4.3 Denoising Configurations
For continuous models like RF, timestamp sampling is of vital importance. Conventionally, RF [29] is trained to predict velocity across timestamps uniformly in $[0,1]$. Considering $\boldsymbol{x}_{t}=t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0}$, prior works [9] uncover that the velocity target at intermediate timestamps, i.e., $t\approx 0.5$, is more challenging. In our case, sampling intermediate timestamps more frequently could increase the difficulty of the reconstruction task, thus effectively amplifying $I(V;G_{1})$ and allowing the visual encoder $\boldsymbol{v}_{\theta}$ to effectively acquire useful fine-grained knowledge from $G_{1}$. In this regard, we propose scaled Logit-Normal sampling for timestamps, as shown below:
@@ 4.4 Generation Paradigms
We choose RF as the continuous denoiser, which is modeled in the latent space of pre-trained VAE [19]. The structure is inherited from FLUX.1-dev [22], consisting of $n\times$ Multimodal Diffusion Transformer (MM-DiT) [33, 9] blocks and $2n\times$ single-stream DiT (Single-DiT) blocks, as shown in Fig. 3 (b). By default, we set $n=2$, which is very efficient with $\sim 1/10$ parameters of the original FLUX.1-dev denoiser. Similar to DiT [33, 9], the condition of visual tokens ([CLS] of $\boldsymbol{v}_{\theta}$) is introduced through the modulation mechanism via adaptive layernorm [33, 1]. The learning objective is the regression of flow matching in Eq. (3).
@@ 5.1 Experimental Setup
For continuous generative models, we choose RF, whose structure is similar to FLUX.1-dev [22], but with only 2 MM-DiT and 4 Single-DiT blocks ($\sim 10\%$ of the parameters). The discrete denoiser is parameterized by a 6-layer Perceiver to predict the masked tokens indexed VQ-GAN’s codebook [8]. Similar to [50], the mask ratio is randomly sampled from 50% to 90%. For both generative models, we only take the [CLS] token of CLIP ViT as the conditional input while dropping other local tokens to prevent information leakage. We choose the scale factor in Eq. (6) as 1 by default.
@@ 5.2 Comparative Results
MMVP- MLLM [42] NaturalBench [24] CV-Bench 2D [41] CV-Bench 3D [41] POPE [25] SciQA- IMG [30] Hallusion Avg. [13]
@@
CLIP has inherently strong global semantics in classification-based tasks. To explore how fine-grained enhancements affect this ability, we evaluate zero-shot classification on datasets like ImageNet-1K [6], CIFAR100 [21], Stanford Cars [20], and SUN397 [49] and zero-shot cross-modal retrieval tasks on Flick30k [53] and COCO [4]. Table 3 reveals that the performance difference is minimal ($<0.3\%$) across various settings, which means that our method could enhance CLIP’s fine-grained understanding without forgetting its global semantics.
@@ 5.4 Further Analysis
In Table 1, we observe that the improvement on SigLIP is relatively smaller compared to OpenAICLIP and MetaCLIP. Specifically, the performance gain over the original SigLIP is $\sim 3.7\%$, less than that for others, i.e., $>10\%$. Unlike the other two backbones [34, 51], SigLIP [55] does not explicitly train a distinct class token. In practice, we extract the pooler_output of SigLIP as the condition for the denoiser, which is obtained by aggregating all local tokens through attention and linear layers. We attribute the relatively small improvement on SigLIP to the indirect leakage of local information through the pooler_output, which hinders the enhancement of $I(V;G_{1})$. This is consistent with the discussion in Sec. 4.2 and the results in Fig. 5.
@@ Appendix A Relationship with Prior Works
DIVA is a pioneer work and proposes to enhance visual representations of CLIP ViT through diffusion feedback. It independently enhances CLIP ViT’s visual representations with the guidance of pre-trained stable diffusion [35]. Similar to DIVA, our work focuses on enhancing CLIP ViT’s internal visual representations. The enhanced CLIP itself could be a more competent vision-language model with better image-text retrieval performance. Furthermore, the enhanced CLIP ViT serves as a plug-and-play module and could be seamlessly plugged into MLLMs. When using the same training recipes but with the enhanced vision encoder, MLLMs could be more capable on several vision-centric benchmarks, with better fine-grained perception of visual details and overcoming visual shortcomings brought about by the original CLIP.
@@ Appendix B More Training Details
Our training process consists of two stages, each involving one epoch on the CC3M [36] dataset. We choose AdamW as the optimizer, with a learning rate of 1e-4 and 1e-5 for Stage-1 and Stage-2, respectively. At Stage-2, we optimize the visual encoder using LoRA with a rank of 16. We train the model on 8 GPUs with a per-device batch size of 16, and the gradient accumulation steps are set as 2, resulting in a global batch size of 256. We plug LoRA to CLIP ViT, with a rank of 16, and an $\alpha$ of 16, and we also employ dropout with a ratio of 0.1 within LoRA.
@@ Appendix D More Experimental Results
Figure 10: Qualitative results of CLIP on MMVP-VLM benchmark. The enhanced CLIP overcomes original visual shortcomings in fine-grained details.
@@ Appendix G Algorithms
5:Pre-trained VAE $\texttt{vae}(\cdot)$ to provide latent space for generative modeling.

### v3

@@ 1 Introduction
Key Point #2: How to design denoising configurations to transfer useful information for visual representations? The structure of the denoiser could determine the enhancement effects. Additionally, before training CLIP, it is essential to mitigate irrelevant information. Therefore, we investigate the influence of different sizes of the denoiser and training stages. In this paper, we propose GenHancer, a two-stage post-training method for visual enhancements. In the first stage, we pre-train the projector and denoiser while freezing the ViT, learning basic reconstruction abilities and mitigating irrelevant information. In the second stage, we fine-tune CLIP ViT to enhance its fine-grained visual representations. Meanwhile, we empirically found that a lightweight denoiser is sufficient to achieve remarkable results, which is more efficient yet stronger, as in Fig. 2.
@@ 2 Related Works
Currently, MLLMs [15, 28] predominantly employ CLIP [41, 47] for visual encoding. Tong et al[53] identified several failure patterns in CLIP, which hinder the fine-grained visual understanding. To overcome this issue, early efforts [53, 52, 20] employed an ensemble of visual experts to combat the visual shortcomings. More recently, ROSS [57] leverages intrinsic visual activations and incorporates a self-supervised visual reconstruction loss during training MLLMs. Complementarily, DIVA [58] proposes to enhance CLIP’s fine-grained abilities through diffusion feedback. Similar to [58], we independently enhance CLIP’s internal representations, which not only strengthen CLIP as a vision-language retriever but also enable the enhanced CLIP to be seamlessly integrated into MLLMs in a plug-and-play manner for better fine-grained vision-centric performance.
@@
Early works [51, 34, 44] utilize generative models as data augmenters [36, 37, 49]. Another line of works [59, 18, 5] leverages self-supervised reconstruction tasks with diffusion models, which helps models grasp visual details and learn fine-grained representations. Similarly, DIVA [58] takes CLIP’s features as conditional inputs to the diffusion model [42], addressing its visual shortcomings through reconstruction. In summary, prior arts predominantly rely on diffusion models [11, 42], whereas we apply our method to both continuous and discrete generative models.
@@
Canonical evaluations of MLLMs focus on fundamental multimodal Q&A capabilities across various domains, e.g., general perception and cognition [10], text and characters [45], scientific fields [33], and potential hallucinations [27, 13] in MLLMs. However, these benchmarks could not effectively assess a model’s fine-grained [14, 15] visual perception abilities, such as object color, quantity, orientation, and viewpoint. To solve this issue, Tong et al[53] systematically explore the failure modes of CLIP and propose a challenging MMVP benchmark with 9 visual patterns. CV-Bench [52] further expands with 2,600 vision-centric VQA questions, covering dimensions like spatial relationships, count, depth, and distance of both 2D and 3D domains. Besides, NaturalBench [26] curates natural adversarial samples that are easy for humans but MLLMs struggle with. In this paper, we employ these vision-centric benchmarks to comprehensively evaluate models’ fine-grained visual abilities.
@@ 3 Preliminaries of Generative Models
$\mathcal{L}_{\text{CE}}=\mathbb{E}_{\boldsymbol{x}\sim p_{\text{data}}}-\log\prod_{i=1}^{L}\boldsymbol{p}_{\phi}(s_{i}|s_{<i}),\vskip-7.0pt$ (2)
@@ 4.3 Denoising Configurations
For continuous models like RF, timestamp sampling is of vital importance. Conventionally, RF [32] is trained to predict velocity across timestamps uniformly in [0,1][0,1][ 0 , 1 ]. Considering $\boldsymbol{x}_{t}=t\boldsymbol{x}_{1}+(1-t)\boldsymbol{x}_{0}$, prior works [9] uncover that the velocity target at intermediate timestamps, i.e., $t\approx 0.5$, is more challenging. In our case, sampling intermediate timestamps more frequently could increase the difficulty of the reconstruction task, thus effectively amplifying $I(V;G_{1})$ and allowing the visual encoder $\boldsymbol{v}_{\theta}$ to effectively acquire useful fine-grained knowledge from $G_{1}$. In this regard, we propose scaled Logit-Normal sampling for timestamps, as shown below:
@@ 4.4 Generation Paradigms
We choose RF as the continuous denoiser, which is modeled in the latent space of pre-trained VAE [21]. The structure is inherited from FLUX.1-dev [24], consisting of $n\times$ Multimodal Diffusion Transformer (MM-DiT) [40, 9] blocks and $2n\times$ single-stream DiT (Single-DiT) blocks, as shown in Fig. 3 (b). By default, we set $n=2$, which is very efficient with 1/10$\sim$ 1/10∼ 1 / 10 parameters of the original FLUX.1-dev denoiser. Similar to DiT [40, 9], the condition of visual tokens ([CLS] of $\boldsymbol{v}_{\theta}$) is introduced through the modulation mechanism via adaptive layernorm [40, 1]. The learning objective is the regression of flow matching in Eq. (3).
@@ 5.1 Experimental Setup
[…]
@@ 5.2 Comparative Results
MMVP-
MLLM [53]
NaturalBench [26] CV-Bench 2D [52]
CV-Bench
3D [52]
POPE [27]
SciQA-
IMG [33]
Hallusion
Avg. [13]
@@
[…]
@@ 5.4 Further Analysis
[…]
@@ Appendix A Relationship with Prior Works
DIVA is a pioneering work and proposes to enhance visual representations of CLIP ViT through diffusion feedback. It independently enhances CLIP ViT’s visual representations with the guidance of pre-trained stable diffusion [42]. Similar to DIVA, our work focuses on enhancing CLIP ViT’s internal visual representations. The enhanced CLIP itself could be a more competent vision-language model with better image-text retrieval performance. Furthermore, the enhanced CLIP ViT serves as a plug-and-play module and could be seamlessly plugged into MLLMs. When using the same training recipes but with the enhanced vision encoder, MLLMs could be more capable on several vision-centric benchmarks, with better fine-grained perception of visual details and overcoming visual shortcomings brought about by the original CLIP.
@@ Appendix B More Training Details
Default training settings.
Our training process consists of two stages, each involving one epoch on the CC3M [43] dataset. We choose AdamW as the optimizer, with a learning rate of 1e-4 and 1e-5 for Stage-1 and Stage-2, respectively. At Stage-2, we optimize the visual encoder using LoRA with a rank of 16. We train the model on 8 GPUs with a per-device batch size of 16, and the gradient accumulation steps are set as 2, resulting in a global batch size of 256. We plug LoRA to CLIP ViT, with a rank of 16, and an $\alpha$ of 16. Additionally, we employ dropout with a ratio of 0.1 within LoRA.
Detailed settings in Fig. 1 of main manuscript.
[…]
@@
Detailed settings in Table 7 of main manuscript.
The pretrained FLUX is also trained under the same setting, i.e., two-stage training and only the [CLS] serves as the condition. Without these proposed keypoints, pretrained denoiser also fails to gain desirable results, e.g., 32.9$\to$22.2 on MMVP, further indicating the generality of our method.
@@ Appendix D More Experimental Results
Figure 9: Qualitative results of CLIP. The enhanced CLIP overcomes the original visual shortcomings in fine-grained details.
@@ Appendix G Algorithms
[…]